# GSAlign: Geometric and Semantic Alignment Network for Aerial-Ground Person Re-Identification

### v1

@@ 4.3 Ablation Study
To assess the individual contributions of each component in the proposed GSAlign framework, we conduct a comprehensive ablation study on the four testing protocols of the CARGO dataset. As shown in Table 4, we start with a strong transformer-based baseline [8] and progressively integrate the Learnable Thin Plate Spline (LTPS) transformation module and the Dynamic Alignment Module (DAM). Compared to the baseline, integrating LTPS brings consistent performance improvements across most settings. Specifically, on the most challenging AG protocol, LTPS improves Rank-1 from 48.12% to 64.89%, mAP from 42.76% to 61.08%, and mINP from 29.95% to 50.54%. These results confirm the effectiveness of progressive geometric alignment in compensating for severe cross-view deformations. Interestingly, the performance gain on AA and GG is more modest, indicating that LTPS is particularly beneficial when strong viewpoint discrepancies exist. Adding the DAM module on top of LTPS yields further improvements, especially in terms of mAP and mINP. Under the full matching scenario (Protocol 1: ALL), the full model achieves the best performance with 65.06% Rank-1, 57.95% mAP, and 44.97% mINP. Notably, the largest relative gain comes from the AG setting, where DAM increases mINP from 50.54% to 52.81%. This demonstrates that DAM effectively suppresses noisy and occluded regions by leveraging visibility-aware semantic masking, thus enhancing cross-view alignment. Overall, the ablation study verifies the complementary effects of LTPS and DAM. While LTPS resolves geometric distortions in a layer-wise manner, DAM provides semantic-level refinement by selectively filtering out unreliable features. Their combination allows GSAlign to maintain robust performance across diverse matching protocols, especially under severe viewpoint and occlusion conditions.
@@
Different variants of DAM. We compare three variants of the Dynamic Alignment Module (DAM): Inner-Batch, Memory Bank, and Classification Matrix. As shown in Table 7, the Inner-Batch variant achieves the best overall performance. It dynamically generates visibility-aware masks within each mini-batch, enabling query-driven refinement of peer features. The Memory Bank variant stores identity prototypes updated by momentum, providing semantic context but suffering from stale or mismatched memory. The Classification Matrix variant reuses classifier weights as class centers [38], but these prototypes are biased toward classification objectives. Overall, Inner-Batch masking proves most effective for dynamic alignment in aerial-ground person ReID.

### v2

@@ 4.3 Ablation Study
To assess the individual contributions of each component in the proposed GSAlign framework, we conduct a comprehensive ablation study on the four testing protocols of the CARGO dataset. As shown in Table 4, we start with a strong transformer-based baseline Zhang et al. (2024b) and progressively integrate the Learnable Thin Plate Spline (LTPS) transformation module and the Dynamic Alignment Module (DAM). Compared to the baseline, integrating LTPS brings consistent performance improvements across most settings. Specifically, on the most challenging AG protocol, LTPS improves Rank-1 from 48.12% to 64.89%, mAP from 42.76% to 61.08%, and mINP from 29.95% to 50.54%. These results confirm the effectiveness of progressive geometric alignment in compensating for severe cross-view deformations. Interestingly, the performance gain on AA and GG is more modest, indicating that LTPS is particularly beneficial when strong viewpoint discrepancies exist. Adding the DAM module on top of LTPS yields further improvements, especially in terms of mAP and mINP. Under the full matching scenario (Protocol 1: ALL), the full model achieves the best performance with 65.06% Rank-1, 57.95% mAP, and 44.97% mINP. Notably, the largest relative gain comes from the AG setting, where DAM increases mINP from 50.54% to 52.81%. This demonstrates that DAM effectively suppresses noisy and occluded regions by leveraging visibility-aware semantic masking, thus enhancing cross-view alignment. Overall, the ablation study verifies the complementary effects of LTPS and DAM. While LTPS resolves geometric distortions in a layer-wise manner, DAM provides semantic-level refinement by selectively filtering out unreliable features. Their combination allows GSAlign to maintain robust performance across diverse matching protocols, especially under severe viewpoint and occlusion conditions.
@@
Different variants of DAM. We compare three variants of the Dynamic Alignment Module (DAM): Inner-Batch, Memory Bank, and Classification Matrix. As shown in Table 7, the Inner-Batch variant achieves the best overall performance. It dynamically generates visibility-aware masks within each mini-batch, enabling query-driven refinement of peer features. The Memory Bank variant stores identity prototypes updated by momentum, providing semantic context but suffering from stale or mismatched memory. The Classification Matrix variant reuses classifier weights as class centers Tan et al. (2022), but these prototypes are biased toward classification objectives. Overall, Inner-Batch masking proves most effective for dynamic alignment in aerial-ground person ReID.